# Let Real Images be as a Judger, Spotting Fake Images Synthesized with Generative Models

### v1

@@ Abstract
In the last few years, generative models have shown their powerful capabilities in synthesizing realistic images in both quality and diversity (i.e., facial images, and natural subjects). Unfortunately, the artifact patterns in fake images synthesized by different generative models are inconsistent, leading to the failure of previous research that relied on spotting subtle differences between real and fake. In our preliminary experiments, we find that the artifacts in fake images always change with the development of the generative model, while natural images exhibit stable statistical properties. In this paper, we employ natural traces shared only by real images as an additional predictive target in the detector. Specifically, the natural traces are learned from the wild real images and we introduce extended supervised contrastive learning to bring them closer to real images and further away from fake ones. This motivates the detector to make decisions based on the proximity of images to the natural traces. To conduct a comprehensive experiment, we built a high-quality and diverse dataset that includes generative models comprising 6 GAN and 6 diffusion models, to evaluate the effectiveness in generalizing unknown forgery techniques and robustness in surviving different transformations. Experimental results show that our proposed method gives 96.1% mAP significantly outperforms the baselines. Extensive experiments conducted on the widely recognized platform Midjourney reveal that our proposed method achieves an accuracy exceeding 78.4%, underscoring its practicality for real-world application deployment. The source code and partial self-built dataset are available in supplementary material.
1 Introduction
@@
With the rapid development and maturity of generative models, the increasing proliferation of fake images has attracted widespread attention. Compared to Generative Adversarial Networks (GANs), diffusion models (DMs), as today’s SOTA generative models, exhibit better generation quality Dhariwal and Nichol (2021) and even support powerful text-to-image models such as DALL-E2 Ramesh et al. (2022), Stable Diffusion Rombach et al. (2022). Currently, one can use different types of generative models to create realistic faces or complex scene images and it is foreseeable that more generative models for image synthesis will emerge in the future. Therefore, it is the goal for the community to develop a more practical method to distinguish fake images that are synthesized with unknown forgery techniques as the unseen generative models will emerge inadvertently.
Figure 1: As the specific artifacts disappear, can fake detectors still work? For example, the detector trained on dog images generated by GANs is unable to detect ones generated by diffusion models.
The prior paradigm for fake image detection is to learn artifacts of fake images by capturing the subtle differences between real and fake images  Zhang et al. (2019); Wang et al. (2020). As shown in Figure 1, the detector distinguishes images as real or fake based on the presence of specific artifacts in the image. It is challenging to generalize to other unknown generative models, which may carry unseen or even imperceptible artifacts. Thus, a practical fake image detector should have strong capabilities in tackling fake images manipulated by unknown generative models. Unfortunately, the existing studies are trapped in the endless efforts to spot the artifacts by investigating the subtle differences between real and fake Corvi et al. (2023). The artifacts commonly explored are ad hoc to generative models and not applicable to diverse forgery techniques. As a defender, fake image detection seems to be perpetually behind the generation, only able to regulate obsolete fake images and helpless against the rapidly updating ones. Such a mouse-and-cat game reminds us whether the paradigm is trapped by exploring the subtle differences between real and fake, which is an endless effort as new forgery methods appear rapidly and the exploited artifacts for forensics will be invalid soon.
We argue that the challenges in the traditional paradigm arise from classifiers’ tendency to easily capture perceptible artifacts in training data. As these artifacts evolve or even vanish with generative model iterations, classifiers fail to detect new fake images. Merely investigating the flaws in fake images proves insufficient, given the constant emergence of unknown fake images that defenders have not seen. Instead, we consider a novel insight that the focus for fake image detection should be on the inherent similarities within real images, as opposed to the differences between real and fake. Our novel insight is motivated by the premise that real-world images possess an intrinsic naturalness that remains stable, whereas fake images continually evolve with technological innovation. We conjecture that there should be some shared features in real images, akin to that in fake images, which we term the natural trace.
Despite the stable statistical properties of natural images, precisely defining and capturing the shared feature of natural traces remains a significant challenge. In this work, we introduce Natural Trace Forensics (NTF) and adopt an alternative strategy, substituting the shared feature with homogeneous features, which are universally present in natural images. We develop a self-supervised feature mapping process to decouple real image representations, capturing stable homogeneous features accurately. Then, we integrate these features as an extra predictive target in parallel with binary classification, prompting the detector to aggregate these features closer to real images and separate them from fake ones. Guided by this strategy, the detector reduces its reliance on specific artifact patterns and gains the capability to detect images synthesized by unknown generative models.
@@
To better evaluate whether our proposed method can be well generalized to other unknown generative models, we build a generated image dataset consisting of 12 SOTA generative models, including 6 GANs and 6 DMs. Our dataset covers a variety of categories, such as faces and scenes, to evaluate detectors’ capability across various types of fake images. In addition, we also evaluate the performance of our method in identifying images generated by Midjourney, which is currently a popular commercial tool for text-to-image generation. Experimental results show that our proposed method could discriminate GAN-based, DM-based, and Multi-step fake images (synthesized with at least two different generated models) in high confidence with an average accuracy of more than 96.1% and is sufficiently robust to various image perturbation transformations.
@@
  • We are the first to propose training fake image detection with natural traces, rather than merely a mix of real and fake images; our results also confirm its ability to detect different generative models and its potential as a universal judge for distinguishing fake images.
  • To conduct a comprehensive evaluation, we build a dataset including GAN-based, DM-based, and Multi-step manipulation for generating fake images. For the first time, we generate Multi-step fake images by employing multiple synthesis methods. This self-built dataset will be released to the community to advance the community to develop more effective and practical methods for distinguishing unseen forgery techniques.
  • Experimental results show the effectiveness of our proposed method in tackling fake images generated by SOTA GANs and diffusion models, giving an average precision of more than 96.18%, significantly outperforming the baselines. Additionally, our method is robust against the common image transformations and identifies the undistinguishable Midjourney in high confidence.
2 Related Work
2.1 Fake Images Synthesis
In recent years, generative models represented by GANs and diffusion models have been evolving towards maturity. GANs train a generator and a discriminator to compete with each other until the generator can produce realistic images, while DMs utilize diffusion and denoising Markov chains for training, surpassing GANs in training stability, diversity, and quality of generation Mazé and Ahmed (2023); Stypułkowski et al. (2024). This marks an evolution from producing or manipulating facial appearances, commonly termed DeepFakes Juefei-Xu et al. (2022), to generating contextually relevant text or image prompts. What’s more, with the emergence of new generative models, the artifacts of fake images are varied and more imperceptible, which also leads to the inability to detect fake images with low-level artifacts. It is foreseeable that the vast majority of artifact-based detection will fail when the fake image artifacts vanish. This progression has escalated public apprehension regarding the challenges in identifying fake images. In this paper, we are dedicated to developing a practical method for detecting fake images that are manipulated by unknown generative models.
2.2 Fake Images Detection
@@
Explicit-based. Some researchers noticed that they often contained specific artifacts or unnatural patternsZhang et al. (2019); Liu et al. (2023). Additionally, several studies focused on exploring GAN-based artifacts in the frequency domainFrank et al. (2020) and the failure to accurately re-enact certain biological features when generating fake facesHu et al. (2021); Tan et al. (2023b). These findings have motivated researchers to use explicit artifacts to detect fake images through simple classifiers. However, as generative models are continuously updated and iterated, these artifacts become imperceptible or even disappear, making detection methods reliant on explicit artifacts less effective.
Implicit-based. Wang et al.Wang et al. (2020) demonstrated that with appropriate training data and data augmentation, neural networks could detect other GAN-based images, while Tan et al. (2023a) used CNNs to transform images into gradient form to present a broader range of artifacts. These studies indicate that images generated by both GAN and diffusion models possess distinct “fingerprints" different from the real imagesYu et al. (2019); Sha et al. (2023). Nevertheless, there is still insufficient evidence to prove that generative models from different families have the same fingerprints that can be used for universal detection.
In this work, we propose an innovative method: training with real images for fake detection. It’s noteworthy that  Liu et al. (2022a) has similarities with ours, where they detect fake by analyzing noise patterns in real images. However, their method relies on the noise patterns of various devices. Our method uses natural traces extracted from real images for training. Compared to noise patterns, which may vary with different devices, can be complex to analyze, and are susceptible to intentional modification for evasion, natural traces consistently present in real images offer greater robustness.
@@
GAN- ProjectedGAN 2021 FFHQ/LSUN/Landscape 12.0k
@@
3 A Diverse Generated Image Dataset
@@
Our dataset covers two major families: GAN and Diffusion Models, with each fake image synthesized using a single generative model. Particularly, we have developed a novel multi-step fake image generation method, involving collaboration between two or more generative models, to achieve identity swapping or attribute editing between real and fake faces. To ensure diversity, the dataset comprises various categories of generation methods: unconditional generation, image-to-image generation, and prompt-guided generation, as shown in Table 1. Moreover, our carefully selected generative models exhibit fundamental differences in their generations, ensuring extensive representation of the dataset.
GAN-based forgery: We select six representative GANs to generate forgery images including unconditional and image-to-image. The innovation of ProGAN Karras et al. (2017) in introducing progressive training has a significant impact on subsequent research, while StyleGAN2 Karras et al. (2020) refines style control through decoupling image features and generates more realistic images. In terms of architectural improvements, ProjectedGAN Sauer et al. (2021) enhances generator feedback using pre-trained weights, while VQGAN Esser et al. (2021) and Diff-StyleGAN2 Wang et al. (2022) innovatively replace the backbone network with transformer and diffusion processes, respectively, resulting in higher image quality and more stable training processes. SimSwap’sChen et al. (2020) ID injection module achieves breakthroughs in arbitrary face swapping.
DM-based forgery: DM-based forgery surpasses GAN-based in terms of image quality and diversity. Here, we select six different DMs for creating fake images including unconditional, image-to-image, and prompt-guided. DDPMHo et al. (2020), as the initial diffusion model, lays the groundwork, with DDIMSong et al. (2020) and PNDMLiu et al. (2022b) enhancing execution speed and quality. DiffFaceKim et al. (2022) is the first identity-conditioned DDPM that uses diffusion models for face swapping. And LDMRombach et al. (2022), which employs pre-trained self-encoders to map pixels to latent space, combined with context learning in cross-attention layers for prompt-guided image generation, along with Stable Diffusion (StableDM), a popular LDM-based prompt-guided model, represent further advancements.
@@
In summary, our dataset includes high-resolution images ranging from $256\times 256$ to $1024\times 1024$ resolution. What’s more, diverse image categories support a more comprehensive evaluation of fake detection generalization. The dataset will be released to the community to advance this interesting research field.
Figure 2: NTF architecture. We first decouple the feature representation of real images into homogeneous and heterogeneous features. Next, the homogeneous features $z^{hom}$ participate in extended supervised contrastive training with the real and fake image features. Detector classifies real and fake images guided by the target of intra-class aggregation and inter-class separation. Better view in color.
Method ProGAN StyleGAN2 VQGAN ProjectedGAN Diff-StyleGAN2 SimSwap Average
@@
XceptionNet 99.31 84.47 0 31.05 100 99.95 0.10 0 100 99.92 0.15 0 100 99.10 1.80 0 97.89 84.03 0.20 31.75 99.85 90.13 19.74 0 99.51 92.93 3.66 10.47
Wang2020 97.88 85.33 27.05 2.00 86.15 56.33 87.30 0.05 61.21 50.52 98.90 0.05 62.18 50.35 99.20 0.10 87.79 68.45 61.30 1.80 98.36 65.04 69.78 0.05 82.26 62.67 73.92 0.68
Grag2021 100 99.98 0 0.05 100 100 0 0 99.88 91.62 16.75 0.05 99.72 94.62 10.75 0 99.82 94.47 11.05 0 94.12 50.49 98.90 0 98.92 88.53 22.91 0.02
Ojha2023 98.83 99.05 1.45 0.45 83.05 87.00 15.95 10.05 83.56 87.30 16.05 9.35 75.20 78.72 33.20 9.35 64.57 64.95 69.30 0.80 80.30 83.83 23.45 8.90 80.92 83.48 26.57 6.48
Self-Blended* 68.65 66.02 39.81 28.16 65.45 60.19 41.75 37.86 56.88 46.15 76.92 30.77 59.87 50.00 53.85 46.15 58.49 54.85 58.25 32.04 55.33 53.88 33.10 59.14 60.78 55.18 41.78 41.95
CADDM* 51.51 53.36 40.63 52.58 52.29 54.41 46.84 44.34 52.62 54.52 50.65 40.28 54.29 57.63 43.00 41.74 51.26 52.90 44.15 50.03 46.64 36.58 82.90 43.63 52.09 53.31 48.36 44.98
NTF 100 92.53 0 14.85 100 92.53 0 14.95 100 92.55 0 14.90 99.77 92.20 0.65 14.95 99.94 92.65 0.15 14.55 99.77 92.16 0.60 15.10 99.91 92.44 0.23 14.88
Table 2: Intra-family generalization on GANs. Performance of NFT and baselines in spotting 6 GAN-based generated images. These baselines include the classic approach, XcetionNet, Wang2020, and Grag2021, as well as the latest methods, Ojha2023, and CADDM. The method with *** means that the baseline is only evaluated on non-face images. The Average column represents the weighted average of the corresponding metrics. Among those, the best and second-best performances are highlighted in bold and underlined, respectively.
Method DDPM DDIM PNDM LDM StableDM DiffFace Average
@@
XceptionNet 74.69 72.37 5.07 50.19 75.78 73.89 12.36 39.86 89.61 49.75 99.95 0.61 65.75 65.02 8.80 61.16 41.07 40.25 28.47 91.03 59.23 40.77 99.45 19.01 66.21 57.01 14.04 59.03
Wang2020 58.15 50.15 99.65 0.05 60.90 50.15 99.60 0.10 66.35 49.95 99.55 0.05 67.87 50.40 99.20 0 39.81 49.48 99.55 1.50 34.75 49.75 99.79 2.40 54.64 49.98 99.56 0.68
Grag2021 79.92 50.12 99.75 0 83.70 50.25 99.45 0.05 86.46 50.10 99.31 0 99.98 98.78 2.45 0 98.17 63.55 72.90 0 63.19 54.35 91.29 0 85.24 61.19 77.53 0.01
Ojha2023 64.18 67.03 57.40 8.55 71.78 75.30 39.65 9.75 61.66 64.42 62.20 8.95 80.20 83.88 22.85 9.40 55.60 55.95 87.30 0.80 88.41 88.72 21.95 0.62 70.31 72.55 48.58 6.34
Self-Blended* 63.12 64.00 37.50 34.50 71.10 69.45 15.53 47.57 64.15 59.22 48.54 33.01 83.75 75.73 14.56 33.98 56.97 55.83 59.22 29.13 70.59 56.91 27.98 28.42 68.28 63.52 33.89 34.44
CADDM* 53.04 55.15 46.47 43.21 55.76 59.77 35.84 44.62 51.75 52.45 51.92 43.10 52.30 53.92 48.28 43.85 51.58 54.51 38.25 52.46 53.02 52.62 43.92 44.15 52.91 54.74 44.11 45.23
NTF 76.93 68.77 47.15 15.30 91.79 84.05 16.40 15.50 79.40 71.02 42.72 15.10 99.98 92.65 0.05 14.65 99.54 92.10 1.05 14.75 99.31 87.43 0.62 15.45 91.16 82.67 18.00 15.13
@@
Method SimSwap_Style2 SimSwap_VQ SimSwap_LDM DiffFace_Style2 DiffFace_Proj DiffFace_LDM Average
@@
XceptionNet 44.60 50.00 100 0 46.56 50.00 100 0 48.51 50.00 100 0 100 100 0 0 100 100 0 0 37.97 50.00 100 0 62.94 66.67 66.67 0
Wang2020 89.32 55.7 88.60 0 76.07 50.80 98.20 0.20 79.38 50.00 98.4 0 69.89 78.91 90.88 0.05 65.19 50.10 99.40 0.40 54.68 50.00 100 0 72.42 55.92 95.91 0.11
Grag2021 99.76 85.70 28.60 0 97.07 61.90 76.00 0 95.24 55.40 89.20 0 100 99.85 0.66 0 80.78 79.40 41.20 0 53.03 50.00 100 0 87.65 72.04 55.94 0
Ojha2023 88.92 91.60 6.60 10.20 90.29 94.10 2.60 9.20 89.27 93.40 3.00 10.20 88.21 92.12 6.14 9.62 87.21 92.00 3.60 12.40 80.96 83.59 26.56 6.25 87.48 91.14 8.08 9.65
Self-Blended* 62.74 69.23 38.46 23.08 74.53 65.49 30.97 38.05 47.38 48.23 74.34 28.20 74.59 69.91 24.78 35.40 65.24 58.33 43.97 45.15 63.04 58.15 54.11 28.80 64.59 61.56 44.44 33.11
CADDM* 46.84 37.30 83.60 41.46 44.56 50.28 60.25 41.14 48.12 44.66 65.80 44.72 18.38 51.11 81.16 40.71 23.48 50.55 67.22 42.71 53.30 55.56 53.74 41.60 39.11 48.24 68.63 42.06
NTF 99.80 91.80 0.40 16.00 99.99 92.20 3.32 15.40 99.87 92.00 0.20 15.80 99.98 87.82 0 15.68 99.66 91.80 0.80 15.60 84.61 74.30 35.60 15.80 97.32 88.32 6.72 15.74
@@
4 Our Method
In this work, to achieve a universal fake image detection and avoid the classifier’s reliance on specific artifact patterns, we propose a novel method, named Natural Trace Forensics (NTF), which involves training the classifier using the natural traces shared merely by real images as an additional predictive target. Figure 2 overviews the pipeline of NTF. We start by learning natural trace representations from real datasets. Then, under an extended supervised contrastive learning framework, the network is trained to align natural traces closer to real images and further from fake ones. With such constraints, the network is motivated to detect fakes based on the distance between images and natural traces. Next, we elaborate on how to learn the natural traces and apply the extracted traces for detection.
4.1 Natural Trace Representation Learning

### v2

# Transfer Learning of Real Image Features with Soft Contrastive Loss for Fake Image Detection

@@ Abstract
In the last few years, the artifact patterns in fake images synthesized by different generative models have been inconsistent, leading to the failure of previous research that relied on spotting subtle differences between real and fake. In our preliminary experiments, we find that the artifacts in fake images always change with the development of the generative model, while natural images exhibit stable statistical properties. In this paper, we employ natural traces shared only by real images as an additional target for a classifier. Specifically, we introduce a self-supervised feature mapping process for natural trace extraction and develop a transfer learning based on soft contrastive loss to bring them closer to real images and further away from fake ones. This motivates the detector to make decisions based on the proximity of images to the natural traces. To conduct a comprehensive experiment, we built a high-quality and diverse dataset that includes generative models comprising GANs and diffusion models, to evaluate the effectiveness in generalizing unknown forgery techniques and robustness in surviving different transformations. Experimental results show that our proposed method gives 96.2% mAP significantly outperforms the baselines. Extensive experiments conducted on popular commercial platforms reveal that our proposed method achieves an accuracy exceeding 78.4%, underscoring its practicality for real-world application deployment.
Introduction
@@
With the rapid development and maturity of generative models, the increasing proliferation of fake images has attracted widespread attention. Compared to Generative Adversarial Networks (GANs), diffusion models (DMs), as today’s SOTA generative models, exhibit better generation quality (Dhariwal 2021) and even support powerful text-to-image models such as DALL-E2 (Ramesh et al. 2022), Stable Diffusion (Rombach et al. 2022). Currently, one can use different types of generative models to create realistic faces or complex scene images and it is foreseeable that more generative models for image synthesis will emerge in the future. Therefore, it is the goal for the community to develop a more practical method to distinguish fake images that are synthesized with unknown forgery techniques as the unseen generative models will emerge inadvertently.
The prior paradigm for fake image detection is to learn artifacts of fake images by capturing the subtle differences between real and fake images  (Wang et al. 2019, 2020; Huang et al. 2020; Wang et al. 2021). Our empirical research confirms that there is a huge challenge in detecting unknown fake images: the model can easily form a classification manifold on the images of the training set, while new generative fake images and their distinctive artifacts will be randomly scattered in this space, as shown in Figure 1 left. Unfortunately, the existing studies are trapped in the endless efforts to spot the artifacts by investigating the subtle differences between real and fake (Corvi et al. 2023). Since the classifier can easily overfit the fake image artifacts of the training set, unknown fake images will be scattered outside the manifold like real images.
Figure 1: TSNE visualization. Left: The detector is easily overfitted to fake images in the training set; Right: Training with natural traces can generalize to unknown fake images.
In this paper, we propose a transfer learning method based on natural trace features for fake image detection. We propose a novel perspective on fake image detection, emphasizing the inherent similarities among real images rather than the differences between real and fake. Our insight is grounded in the premise that images from the real world possess a stable intrinsic naturalness. We posit that real images share common characteristics, akin to those found in fake images, which we refer to as "natural traces." We introduce natural trace forensics (NTF) and adopt a substitution strategy to replace shared features with homogeneous features. Specifically, we exploit a self-supervised feature mapping process to decouple heterogeneous features of real images, allowing for the capture of stable homogeneous features. We then develop a transfer learning for homogeneous features based on a soft contrastive loss, which simultaneously solved the problem that similar clusters formed by self-supervised contrastive learning are still scattered with large distances and outliers under the constraints of supervised contrastive learning do not converge. We jointly optimized transfer training and binary classification to force the classifier to aggregate homogeneous features close to real images and constrain them away from fake images. In Figure 1 right, the classifier reduces its dependence on specific artifact patterns and gains the ability to detect images synthesized by unknown generative models with natural traces.
@@
To better evaluate whether our proposed method can be well generalized to other unknown generative models, we build a generated image dataset consisting of 12 SOTA generative models, including 6 GANs and 6 DMs. Our dataset covers a variety of categories, such as faces and scenes, to evaluate detectors’ capability across various types of fake images. In addition, we also evaluate the performance of our method in identifying images generated by Midjourney [^1] https://www.midjourney.com/ and Kolors [^2] https://klingai.kuaishou.com/, which is currently a popular commercial tool for text-to-image generation. Experimental results show that our proposed method could discriminate GAN-based, DM-based, and Multi-step fake images (synthesized with at least two different generated models) in high confidence with an average accuracy of more than 96.1% and is sufficiently robust to various image perturbation transformations.
@@
  • We are the first to propose soft contrastive transfer learning, which utilizes the disentangled shared features of real images for fake image detection. Our results also demonstrate its effectiveness in detecting previously unseen generative fake images, highlighting its potential as a versatile solution in this domain.
  • To conduct a comprehensive evaluation, we build a dataset including GAN-based, DM-based, and Multi-step manipulation for generating fake images. For the first time, we generate Multi-step fake images by employing multiple synthesis methods.
  • Experimental results show the effectiveness of our proposed method in tackling fake images generated by SOTA GANs and diffusion models, giving an average precision of more than 96.18%, significantly outperforming the baselines.
Related Work
Transfer Learning in Fake Detection
In recent years, the most common method of transfer learning is to fine-tune the pre-trained model on the target dataset. (Suratkar 2023) detect fake videos through the utilization of transfer learning in autoencoders and a hybrid model of CNN and RNN. (Lee and Kim 2022; Ghayoomi 2022) used pre-trained BiLSTM(Zhou et al. 2016) and RoBerta(Liu et al. 2019) for transfer learning on fake news about COVID-19 in Korean and Persian, respectively. Although the fine-tuning method is highly flexible for downstream tasks, it is also prone to catastrophic forgetting. Freezing layers of the pre-trained backbone network and finetuning only lateral fully connected layers is one of the most effective transfer learning methods. (Ranjan 2020) explored the effectiveness and interpretability of freezing convolutional layers and fine-tuning fully connected layers in DeepFake detection, and (Elhassan et al. 2022) studied the transfer of lip-motion-based DeepFake detection methods on 11 models. However, the pre-training of these frozen transfer learning methods is obviously isolated from the target, and the parameter transfer effect is limited. As far as we know, there is still a gap in the exploration of transfer learning in fake image detection. Our method balances the feature transfer and classification by additional soft contrastive loss constraints.
Fake Images Detection
@@
Explicit-based. Some researchers noticed that they often contained specific artifacts or unnatural patterns(Zhang 2019; Liu et al. 2023). Additionally, several studies focused on exploring GAN-based artifacts in the frequency domain(Frank et al. 2020) and the failure to accurately re-enact certain biological features when generating fake faces(Hu 2021; Tan et al. 2023b). These findings have motivated researchers to use explicit artifacts to detect fake images through simple classifiers. However, as generative models are continuously updated and iterated, these artifacts become imperceptible or even disappear, making detection methods reliant on explicit artifacts less effective.
Implicit-based. Wang et al.(Wang et al. 2020) demonstrated that with appropriate training data and data augmentation, neural networks could detect other GAN-based images, while (Tan et al. 2023a) used CNNs to transform images into gradient form to present a broader range of artifacts. These studies indicate that images generated by both GAN and diffusion models possess distinct “fingerprints" different from the real images(Yu, Davis, and Fritz 2019; Sha et al. 2023). Nevertheless, there is still insufficient evidence to prove that generative models from different families have universal fingerprints for detection.
@@
GAN- ProjGAN 2021 FFHQ/LSUN/Landscape 12.0k
@@
A Diverse Generated Image Dataset
@@
Our dataset covers two major families: GANs and diffusion models, with each fake image synthesized using one generative model. Particularly, we have developed a novel multi-step fake image generation method, involving collaboration between two or more generative models, to achieve identity swapping or attribute editing between real and fake faces. To ensure diversity, the dataset comprises various categories of generation methods: unconditional generation, image-to-image generation, and prompt-guided generation, as shown in Table 1. Moreover, our carefully selected generative models exhibit fundamental differences in their generations, ensuring extensive representation of the dataset.
GAN-based forgery: We select six representative GANs to generate forgery images including unconditional and image-to-image. The innovation of ProGAN (Karras et al. 2017) in introducing progressive training has a significant impact on subsequent research, while StyleGAN2 (Karras et al. 2020) refines style control through decoupling image features and generates more realistic images. In terms of architectural improvements, ProjGAN (Sauer et al. 2021) enhances generator feedback using pre-trained weights, while VQGAN (Esser 2021) and Diff-StyleGAN2 (Wang et al. 2022) innovatively replace the backbone network with transformer and diffusion processes, respectively, resulting in higher image quality and more stable training processes. SimSwap’s (ss)(Chen et al. 2020) ID injection module achieves breakthroughs in arbitrary face swapping.
DM-based forgery: DM-based surpasses GAN-based in terms of image quality and diversity. Here, we select six different DMs for creating fake images including unconditional, image-to-image, and prompt-guided. DDPM(Ho 2020), as the initial diffusion model, lays the groundwork, with DDIM(Song 2020) and PNDM(Liu et al. 2022) enhancing execution speed and quality. DiffFace (df)(Kim et al. 2022) is the first identity-conditioned DDPM that uses diffusion models for face swapping. And LDM(Rombach et al. 2022), which employs pre-trained self-encoders to map pixels to latent space, combined with context learning in cross-attention layers for prompt-guided image generation, along with Stable Diffusion (SDM), a popular LDM-based prompt-guided model, represent further advancements.
@@
Figure 2: NTF architecture. We first decouple the feature representation of real images into homogeneous and heterogeneous features. Next, the homogeneous features $z^{hom}$ participate in SCL with the real and fake image features. Detector classifies real and fake images guided by the target of intra-class aggregation and inter-class separation. Better view in color.
Category Method ProGAN StyleGAN2 VQGAN ProjGAN Diff-StyleGAN2 SimSwap Average
@@
Explicit-based SBIs* 68.6 66.0 39.8 28.2 65.5 60.2 41.8 37.9 56.9 46.2 76.9 30.8 59.9 50.0 53.9 46.2 58.5 54.9 58.3 32.0 55.3 53.9 33.1 59.1 60.8 55.2 41.8 41.9
CADDM* 51.5 53.4 40.6 52.6 52.3 54.4 46.8 44.3 52.6 54.5 50.7 40.3 54.3 57.6 43.0 41.7 51.3 52.9 44.2 50.0 46.6 36.6 82.9 43.6 52.1 53.3 48.4 45.0
Implicit-based Xception 99.3 84.5 0 31.1 100 99.9 0.1 0 100 99.9 0.2 0 100 99.1 1.8 0 97.9 84.0 0.2 31.8 99.9 90.1 19.7 0 99.5 92.9 3.7 10.5
Wang2020 97.9 85.3 27.1 2.0 86.2 56.3 87.3 0.1 61.2 50.5 98.9 0.1 62.2 50.4 99.2 0.1 87.8 68.5 61.3 1.8 98.4 65.0 69.8 0.1 82.3 62.7 73.9 0.7
Grag2021 100 99.9 0 0.1 100 100 0 0 99.9 91.6 16.8 0.1 99.7 94.6 10.8 0 99.8 94.5 11.1 0 94.1 50.5 98.9 0 98.9 88.5 22.9 0
Ojha2023 98.8 99.1 1.5 0.5 83.1 87.0 15.9 10.1 83.6 87.3 16.1 9.4 75.2 78.7 33.2 9.4 64.6 64.9 69.3 0.8 80.3 83.8 23.5 8.9 80.9 83.5 26.6 6.5
NTF(Ours) 100 92.5 0 14.9 100 92.5 0 15.0 100 92.6 0 14.9 99.8 92.2 0.7 15.1 99.9 92.7 0.2 14.6 99.8 92.2 0.6 15.1 99.9 92.4 0.2 14.9
Table 2: Intra-family generalization on GANs. Performance of NFT and baselines in spotting 6 GAN-based generated images. These baselines include the classic approach, XcetionNet, Wang2020, and Grag2021, as well as the latest methods, Ojha2023, and CADDM. The method with * means that the baseline is only evaluated on non-face images. The Average column represents the weighted average of the corresponding metrics. Among those, the best and second-best performances are highlighted in bold and underlined, respectively.
Category Method DDPM DDIM PNDM LDM SDM DiffFace Average
@@
Explicit-based SBIs* 63.1 64.0 37.5 34.5 71.1 69.5 15.5 47.6 64.2 59.2 48.5 33.0 83.8 75.7 14.6 34.0 57.0 55.8 59.2 29.1 70.6 56.9 28.0 28.4 68.3 63.5 33.9 34.4
CADDM* 53.0 55.2 46.5 43.2 55.8 59.8 35.8 44.6 51.8 52.5 51.9 43.1 52.3 53.9 48.3 43.9 51.6 54.5 38.3 52.5 53.0 52.6 43.9 44.2 52.9 54.7 44.1 45.2
Implicit-based Xception 74.7 72.3 5.1 50.2 75.8 73.9 12.4 39.9 89.6 49.8 99.9 0.6 65.8 65.0 8.8 61.2 41.1 40.3 28.5 91.0 59.2 40.8 99.5 19.0 66.2 57.0 14.0 59.0
Wang2020 58.2 50.2 99.7 0.1 60.9 50.2 99.6 0.1 66.4 50.0 99.6 0.1 67.9 50.4 99.2 0 39.8 49.5 99.6 1.5 34.8 49.8 99.8 2.4 54.6 50.0 99.6 0.7
Grag2021 79.9 50.1 99.8 0 83.7 50.3 99.5 0.1 86.5 50.1 99.3 0 99.9 98.8 2.5 0 98.2 63.6 72.9 0 63.2 54.4 91.3 0 85.2 61.2 77.5 0
Ojha2023 64.2 67.0 57.4 8.6 71.8 75.3 39.7 9.8 61.7 64.4 62.2 9.0 80.2 83.9 22.9 9.4 55.6 56.0 87.3 0.8 88.4 88.7 22.0 0.6 70.3 72.6 48.6 6.3
NTF(Ours) 76.9 68.8 47.2 15.3 91.8 84.1 16.4 15.5 79.4 71.0 42.7 15.1 99.9 92.7 0.1 14.7 99.5 92.1 1.1 14.8 99.3 87.4 0.6 15.5 91.2 82.7 18.0 15.1
@@
Category Method SimSwap_Style2 SimSwap_VQ SimSwap_LDM DiffFace_Style2 DiffFace_Proj DiffFace_LDM Average
@@
Explicit-based SBIs* 62.7 69.2 38.5 23.1 74.5 65.5 31.1 38.1 47.4 48.2 74.3 28.2 74.6 69.9 24.8 35.4 65.2 58.3 44.0 45.2 63.0 58.2 54.1 28.8 64.6 61.6 44.4 33.1
CADDM* 46.8 37.3 83.6 41.5 44.6 50.3 60.3 41.1 48.1 44.7 65.8 44.7 18.4 51.1 81.2 40.7 23.5 50.6 67.2 42.7 53.3 55.6 53.7 41.6 39.1 48.2 68.6 42.1
Implicit-based Xception 44.6 50.0 100 0 46.6 50.0 100 0 48.5 50.00 100 0 100 100 0 0 100 100 0 0 38.0 50.0 100 0 62.9 66.7 66.7 0
Wang2020 89.3 55.7 88.6 0 76.1 50.8 98.2 0.2 79.4 50.0 98.4 0 69.9 78.9 90.9 0.1 65.2 50.1 99.4 0.4 54.7 50.0 100 0 72.4 55.9 95.9 0.1
Grag2021 99.8 85.7 28.6 0 97.1 61.9 76.0 0 95.2 55.4 89.2 0 100 99.9 0.7 0 80.9 79.4 41.2 0 53.0 50.0 100 0 87.7 72.0 55.9 0
Ojha2023 88.9 91.6 6.6 10.2 90.3 94.1 2.6 9.2 89.3 93.4 3.0 10.2 88.2 92.1 6.1 9.6 87.2 92.0 3.6 12.4 81.0 83.6 26.6 6.3 87.5 91.1 8.1 9.7
NTF(Ours) 99.8 91.8 0.4 16.0 99.9 92.2 3.3 15.4 99.9 92.0 0.2 15.8 99.9 87.8 0 15.7 99.7 91.8 0.8 15.6 84.6 74.3 35.6 15.8 97.3 88.3 6.7 15.7
@@
Our Method
In this work, we propose a novel method, named Natural Trace Forensics (NTF), which involves training the classifier using the natural traces shared merely by real images as an additional predictive target. Figure 2 overviews the pipeline of NTF. We start by learning natural trace representations from real datasets. Then, under a soft contrastive learning (SCL) framework, the network is trained to align natural traces closer to real images and further from fake ones. With such constraints, the network is motivated to detect fakes based on the distance between images and natural traces. Next, we elaborate on how to learn the natural traces and apply the extracted traces for detection.
Natural Trace Representation Learning